# Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs

### v2

@@ 6.1 Instances
Graphs from the Second DIMACS Implementation Challenge [37]. These form the standard benchmark for max-clique algorithms. We complemented the graphs to convert the max-clique instances into SSP instances (DSJC* graphs are not complemented, as these belong to the “coloring” benchmark set). We excluded the c-fat, johnson, and san graphs because the integrality gap is completely closed by linear relaxations, so they do not provide information to our analysis. We also dropped keller5 and the largest and p-hat instances (with 1000 and 1500 vertices) as the SDP relaxations turned out to be too large to be handled. The final collection includes 38 graphs. Statistics about the time required to compute lifting coefficients of linear formulations $N_{+}(\operatorname{NOD}(G,\theta))$ and $N_{+}(\operatorname{NOD}(G,\alpha))$ are reported in Table 2. Recall that their computation can be carried out in parallel, so the impact on the overall time is often negligible.
@@ 6.2 Experiment 1: random graphs
Table 3 collects gap statistics for Erdös–Rényi graphs, where the first column reports the integrality gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$. Tables 4 and 5 report cutting plane iterations, CPU times, and, respectively, the number of added cuts for each class of inequalities. Let us first notice that $\lambda(G)\simeq\theta^{+}(G)$ in all cases but $n=200$, $p\in\{0.1,0.9\}$, evidencing a rather disappointing behavior of $M_{+}(\operatorname{FRAC}(G))$. On the contrary, replacing edge inequalities with a compact collection of clique inequalities in the initial formulation pays off, as $\mu(G,{\cal C})$ is always smaller than $\lambda(G)$. The average portion of additional gap closed by $\mu(G,{\cal C})$ w.r.t. $\lambda(G)$ is $6.7\%$. This is achieved by a few cutting plane iterations, adding about one to three thousand inequalities of class (19), resulting in an average $80\%$ CPU time increase. Interestingly, no inequalities in the class (21) have been detected as violated.
@@
This experiment clearly classifies relaxations strength as a function of edge probability $p$ and graph size $n$. When $p\leq 0.4$, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ turns out to be the strongest relaxation and the sparser the graph, the larger the gap closed by $\mu(G,{C})$ w.r.t. other relaxations, while formulations based on nodal inequalities fail to improve on $\theta^{+}(G)$. However, the advantage of $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ over other relaxations tends to thin out as $n$ increases. The outcome is completely reversed when $p\geq 0.5$, where relaxations based on nodal inequalities are more competitive and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$ is the clear winner. $\nu(G,\alpha)$ closes $45.3\%$ of the gap left by $\lambda(G)$ and $40.0\%$ of the one left by $\mu(G,{C})$, on average. To our knowledge, such an improvement on $\theta^{+}(G)$ has rarely been documented in the literature on SDP approaches, even for smaller graphs. Another nice evidence is that this advantage of $M_{+}(\operatorname{NOD}(G,\alpha))$ is independent of the graph size. Observe that, also in this case, one class of inequalities is predominant, namely, (25). Notice also that some violated inequalities in (28) and (29) have only been detected for the densest graphs $p=0.9$. Very good upper bounds are also achieved by $M_{+}(\operatorname{NOD}(G,\theta))$ when $p\geq 0.7$. In fact, $\nu(G,\theta)$ remarkably improves on $\theta^{+}(G)$ and, unlike $\nu(G,\alpha)$, can be computed in polynomial time. Also, the computational effort required to achieve strong bounds is reasonable. Differently, the weakest relaxation $M_{+}(\operatorname{NOD}(G,\Gamma))$ always returns $\nu(G,\Gamma)=\theta^{+}(G)$, that is, never closes additional gap in our tests. Finally, notice that for highly dense graphs, the gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$ may be smaller than that of the weakest SDP relaxations. However, it is largely improved by the strongest SDP relaxation, which confirms the operator’s power and the PSD constraint’s key role.
@@ 6.3 Experiment 2: DIMACS instances
The DIMACS test set, a standard for max-clique/stable set studies, includes, for the most part, structured graphs from various applications. The breadth of the graph collection supports the relevance of our analysis. To our knowledge, DIMACS graphs with more than 300-400 vertices have never been documented in SDP-based studies [24, 23, 25, 41]. The graphs we have considered are significantly larger than those in case studies applying the $N_{+}(\cdot)$ operator in [21, 12]. In our experience, experimenting with larger graphs is crucial to draw reliable conclusions about the strength of relaxations for the stable set problem, as it typically tends to degrade with graph size. Looking at Table 6, as already noted for random graphs, $\lambda(G)$ does not prove to be better than $\theta^{+}(G)$: some improvement is observed in two cases (DSJC125-1 and MANN_a27), it is negligible in nine cases and null in the remaining ones. Concerning the relaxations we have proposed, $M_{+}(\operatorname{NOD}(G,\alpha))$ shows the most promising results. $\nu(G,\alpha)$ is the best bound in 11 cases while $\mu(G,{\cal C})$ is the winner in as many as 23 cases. Nevertheless, $\nu(G,\alpha)$ achieves the smallest gap on average, thanks to its remarkable strength on selected instances. According to the evidence gathered from Erdös–Rényi graphs, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ is more suited for sparse instances, and $M_{+}(\operatorname{NOD}(G,\alpha))$ turns out to be particularly tight as the density increases. $\mu(G,{\cal C})$ is the best bound for the C and MANN collections and the sparsest among p_hat graphs. On the other hand, $\nu(G,\alpha)$ is very strong for DSJC500-5, p_hat300-1, p_hat500-1, p_hat700-1, sanr400_0.5, where it closes more than $50\%$ of the gap left by $\theta^{+}(G)$ and $\lambda(G)$ (about $74\%$ for p_hat-1 graphs). The computational price for this is reported in Table 7: computing $\mu(G,\alpha)$ is about 27 times slower than computing $\theta^{+}(G)$ except p_hat700-1 where it is 185 times slower. But this experiment also shows that the tightness of $M_{+}(\operatorname{NOD}(G,\alpha))$ is less sensitive to the graph size with respect to the other relaxations. This results in an interesting outcome for the hard brock800 collection that, with a density of $35\%$, one expects to be favorable to $M_{+}(\operatorname{FRAC}(G))$ and $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$. $\nu(G,\alpha)$ is the only bound stronger than $\theta^{+}(G)$ on these graphs, and the improvement is significant. Table 7 shows that computing $\nu(G,\alpha)$ is about four times slower than computing $\theta^{+}(G)$ (including the time to build the LP relaxation), an acceptable price for such a valuable result. Table 8 also reveals the key contribution to closing the gap given by inequalities (25).
@@
Also in this experiment, the polynomial bound $\nu(G,\theta)$ improves on $\theta^{+}(G)$ in six cases by a meaningful amount (namely, DSJC125-5, MANN_a27, p_hat300-1, p_hat500-1, p_hat700-1). The comparison between $\nu(G,\theta)$ and $\nu(G,\alpha)$ reveals that strengthening the lifting coefficient pays off, as in eleven cases, $\nu(G,\alpha)$ is significantly smaller than $\nu(G,\theta)$, in other eight cases the improvement is negligible while no improvement is observed in the remaining sixteen cases. On the other hand, $\nu(G,\Gamma)$ has never enhanced $\theta^{+}(G)$ even if a few violated cuts have occasionally been detected. Looking at Tables 7 and 8, one can observe that relevant improvements on $\theta^{+}(G)$ are achieved by adding several inequalities in a few iterations of the cutting plane algorithm. Similarly to what is observed for random graphs, inequalities (21) are only added in a few special graphs, namely the extremely sparse MANN graphs. Differently, classes of inequalities (25) and (26) are relevant to describe $M_{+}(\operatorname{NOD}(G,\theta))$, $M_{+}(\operatorname{NOD}(G,\alpha))$, while no violated members of classes (28) and (29) have ever been found.
@@
Overall, the computational analysis shows that the new SDP relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$, may be a viable option to compute upper bounds to $\alpha(G)$ which are remarkably stronger than $\theta(G)$ at a reasonable computational price. Interestingly, some experimental evidence complements the theoretical insights illustrated in Section 4.3, where $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ looks to be likely to achieve the tightest bounds than $M_{+}(\operatorname{NOD}(G,\alpha))$.
7 Conclusions
We have presented a way to improve the classical Lovász $\theta$ bound for the stability number of a graph by applying the Lovász-Schrijver $N_{+}(\cdot)$ lift-and-project operator to tailored LP relaxations. Unlike previous approaches, some of the latter may not be constructed in polynomial time. In the case of clique inequalities, we have to handle exponentially many inequalities (with an associated NP-hard separation problem), while building strong nodal inequalities requires solving the SSP on specific subgraphs. We show that these expedients, if properly handled, are indeed helpful in letting the resulting SDP relaxations improve the $\theta$ bound without excessive extra effort. This study reveals that the Lovász and Schrijver operators can be of practical interest besides being a powerful theoretical tool. Finally, an interesting research direction consists of finding classes of graphs where ${N_{+}(\operatorname{NOD}(G,\theta))}=\operatorname{STAB}(G)$, as for them the SSP would be polynomially solvable.
Acknowledgements
Conflict of interest
The authors declare that they have no conflict of interest.
@@ Acknowledgements
Table 6: Lift-and-project bounds on DIMACS instances. The column $d(\%)$ report the density of each graph.
@@
Name $d$(%) Time # iter Time # iter Time # iter Time # iter Time # iter Time

### v3

@@ 6.1 Instances
Graphs from the Second DIMACS Implementation Challenge [37] form the standard benchmark for max-clique algorithms. We complemented the graphs to convert the max-clique instances into SSP instances (DSJC* graphs are not complemented, as these belong to the “coloring” benchmark set). We excluded the c-fat, johnson, and san graphs because the integrality gap is completely closed by linear relaxations, so they do not provide information to our analysis. We also dropped keller5 and the largest p-hat instances (with 1000 and 1500 vertices) as the SDP relaxations turned out to be too large to be handled. The final collection includes 38 graphs. Statistics about the time required to compute lifting coefficients of linear formulations $N_{+}(\operatorname{NOD}(G,\theta))$ and $N_{+}(\operatorname{NOD}(G,\alpha))$ are reported in Table 2. Recall that their computation can be carried out in parallel, so the impact on the overall time is often negligible.
@@ 6.2 Experiment 1: random graphs
Table 3 collects gap statistics for Erdös–Rényi graphs, where the first column reports the integrality gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$. Tables 4 and 5 report cutting plane iterations and CPU times, and the number of added cuts for each class of inequalities, respectively.
@@
Let us first notice that $\lambda(G)\simeq\theta^{+}(G)$ in all cases but $n=200$, $p\in\{0.1,0.9\}$, evidencing a rather disappointing behavior of $M_{+}(\operatorname{FRAC}(G))$. On the contrary, replacing edge inequalities with a compact collection of clique inequalities in the initial formulation pays off, as $\mu(G,{\cal C})$ is always smaller than $\lambda(G)$. The average portion of additional gap closed by $\mu(G,{\cal C})$ w.r.t. $\lambda(G)$ is $6.7\%$. This is achieved by a few cutting plane iterations, adding about one to three thousand inequalities of class (13b), resulting in an average $80\%$ CPU time increase. Interestingly, no inequalities in the class (13d) have been detected as violated.
@@
This experiment clearly classifies relaxations strength as a function of edge probability $p$ and graph size $n$. When $p\leq 0.4$, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ turns out to be the strongest relaxation and the sparser the graph, the larger the gap closed by $\mu(G,{C})$ w.r.t. other relaxations, while formulations based on nodal inequalities fail to improve on $\theta^{+}(G)$. However, the advantage of $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ over other relaxations tends to thin out as $n$ increases. The outcome is completely reversed when $p\geq 0.5$, where relaxations based on nodal inequalities are more competitive and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$ is the clear winner. $\nu(G,\alpha)$ closes $45.3\%$ of the gap left by $\lambda(G)$ and $40.0\%$ of the one left by $\mu(G,{\cal C})$, on average. To our knowledge, such an improvement on $\theta^{+}(G)$ has rarely been documented in the literature on SDP approaches, even for smaller graphs. Another nice evidence is that this advantage of $M_{+}(\operatorname{NOD}(G,\alpha))$ is independent of the graph size. Observe that, also in this case, one class of inequalities is predominant, namely, (15b). Notice also that some violated inequalities in (16b) and (16c) have only been detected for the densest graphs $p=0.9$. Very good upper bounds are also achieved by $M_{+}(\operatorname{NOD}(G,\theta))$ when $p\geq 0.7$. In fact, $\nu(G,\theta)$ remarkably improves on $\theta^{+}(G)$ and, unlike $\nu(G,\alpha)$, can be computed in polynomial time. Also, the computational effort required to achieve strong bounds is reasonable. Differently, the weakest relaxation $M_{+}(\operatorname{NOD}(G,\Gamma))$ always returns $\nu(G,\Gamma)=\theta^{+}(G)$, that is, never closes additional gap in our tests. Finally, notice that for highly dense graphs, the gap of the LP relaxation $\operatorname{NOD}(G,\alpha)$ may be smaller than that of the weakest SDP relaxations. However, it is largely improved by the strongest SDP relaxation, which confirms the operator’s power and the PSD constraint’s key role.
@@ 6.3 Experiment 2: DIMACS instances
The DIMACS test set, a standard for max-clique/stable set studies, includes, for the most part, structured graphs from various applications. The breadth of the graph collection supports the relevance of our analysis. To our knowledge, DIMACS graphs with more than 300-400 vertices have never been documented in SDP-based studies [23, 24, 25, 41]. The graphs we have considered are significantly larger than those in case studies applying the $N_{+}(\cdot)$ operator in [12, 21]. In our experience, experimenting with larger graphs is crucial to draw reliable conclusions about the strength of relaxations for the stable set problem, as it typically tends to degrade with graph size. Looking at Table 6, as already noted for random graphs, $\lambda(G)$ does not prove to be better than $\theta^{+}(G)$: some improvement is observed in two cases (DSJC125-1 and MANN_a27), it is negligible in nine cases and null in the remaining ones. Concerning the relaxations we have proposed, $M_{+}(\operatorname{NOD}(G,\alpha))$ shows the most promising results. $\nu(G,\alpha)$ is the best bound in 11 cases while $\mu(G,{\cal C})$ is the winner in as many as 23 cases. Nevertheless, $\nu(G,\alpha)$ achieves the smallest gap on average, thanks to its remarkable strength on selected instances. According to the evidence gathered from Erdös–Rényi graphs, $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ is more suited for sparse instances, and $M_{+}(\operatorname{NOD}(G,\alpha))$ turns out to be particularly tight as the density increases. $\mu(G,{\cal C})$ is the best bound for the C and MANN collections and the sparsest among p_hat graphs. On the other hand, $\nu(G,\alpha)$ is very strong for DSJC500-5, p_hat300-1, p_hat500-1, p_hat700-1, sanr400_0.5, where it closes more than $50\%$ of the gap left by $\theta^{+}(G)$ and $\lambda(G)$ (about $74\%$ for p_hat-1 graphs). The computational price for this is reported in Table 7: computing $\mu(G,\alpha)$ is about 27 times slower than computing $\theta^{+}(G)$ except p_hat700-1 where it is 185 times slower. But this experiment also shows that the tightness of $M_{+}(\operatorname{NOD}(G,\alpha))$ is less sensitive to the graph size with respect to the other relaxations. This results in an interesting outcome for the hard brock800 collection that, with a density of $35\%$, one expects to be favorable to $M_{+}(\operatorname{FRAC}(G))$ and $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$. $\nu(G,\alpha)$ is the only bound stronger than $\theta^{+}(G)$ on these graphs, and the improvement is significant. Table 7 shows that computing $\nu(G,\alpha)$ is about four times slower than computing $\theta^{+}(G)$ (including the time to build the LP relaxation), an acceptable price for such a valuable result. Table 8 also reveals the key contribution to closing the gap given by inequalities (15b).
@@
Also in this experiment, the polynomial bound $\nu(G,\theta)$ improves on $\theta^{+}(G)$ in six cases by a meaningful amount (namely, DSJC125-5, MANN_a27, p_hat300-1, p_hat500-1, p_hat700-1). The comparison between $\nu(G,\theta)$ and $\nu(G,\alpha)$ reveals that strengthening the lifting coefficient pays off, as in eleven cases, $\nu(G,\alpha)$ is significantly smaller than $\nu(G,\theta)$, in other eight cases the improvement is negligible while no improvement is observed in the remaining sixteen cases. On the other hand, $\nu(G,\Gamma)$ has never enhanced $\theta^{+}(G)$ even if a few violated cuts have occasionally been detected.
Looking at Tables 7 and 8, one can observe that relevant improvements on $\theta^{+}(G)$ are achieved by adding several inequalities in a few iterations of the cutting plane algorithm. Similarly to what is observed for random graphs, inequalities (13d) are only added in a few special graphs, namely the extremely sparse MANN graphs. Differently, classes of inequalities (15b) and (15c) are relevant to describe $M_{+}(\operatorname{NOD}(G,\theta))$, $M_{+}(\operatorname{NOD}(G,\alpha))$, while no violated members of classes (16b) and (16c) have ever been found. Interestingly, our numerical results indicate that inequalities generated during the lifting step by multiplying by $x_{i}\geq 0$ are more frequently violated than those obtained by multiplying by $1-x_{i}\geq 0$. This trend is consistent across all reported relaxations, with only a few exceptions.
@@
Overall, the computational analysis shows that the new SDP relaxations $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and, mainly, $M_{+}(\operatorname{NOD}(G,\alpha))$, may be a viable option to compute upper bounds on $\alpha(G)$ which are remarkably stronger than $\theta(G)$ at a reasonable computational price. Interestingly, some experimental evidence complements the theoretical insights illustrated in Section 4.3, where $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ looks to be likely to achieve tighter bounds than $M_{+}(\operatorname{NOD}(G,\alpha))$.
@@
Table 6: Lift-and-project bounds on DIMACS instances. The column $d(\%)$ reports the density of each graph.
@@
Name dd(%) Time # iter Time # iter Time # iter Time # iter Time # iter Time
@@
6.4 Sensitivity to the choice of $\mathcal{C}$
To enhance the results of our computational analysis, we investigate whether refining the edge clique cover $\mathcal{C}$ obtained from Algorithm 1 can lead to a significant improvement in the bound $\mu(G,\mathcal{C})$, while also considering the associated computational overhead.
To this end, we consider the edge clique cover $\mathcal{C_{\rm cut}}\in\mathcal{Q}$ proposed by Letchford et al. [40]. The collection $\mathcal{C_{\rm cut}}$ is obtained by adding to $\mathcal{C}$ all cliques identified through a cutting plane algorithm that employs the separation heuristic of Marzi et al. [45]. Since $\mathcal{C}\subseteq\mathcal{C_{\rm cut}}$, $\operatorname{QSTAB}(G,\mathcal{C_{\rm cut}})$ provides a tighter upper bound on $\alpha(G)$ than that of $\operatorname{QSTAB}(G,\mathcal{C})$}. Marzi et al. [45] and Letchford et al. [40] also show that this upper bound is very close to the bound provided by $\operatorname{QSTAB}(G)$}. However, this improvement may require a large number of clique inequalities.
Tables 9 and 10 report the results of this sensitivity analysis on a subset of the previously tested DIMACS instances. We selected graphs of varying densities for which we expect that $\mathcal{C_{\rm cut}}$ has the potential to overturn the conclusion drawn from the previous experiment. As anticipated, the percentage gaps in Table 9 show an improvement of $\mu(G,\mathcal{C_{\rm cut}})$ over $\mu(G,\mathcal{C})$. This gain, however, generally comes at a substantial additional cost in terms of cutting-plane iterations and CPU time (up to a factor of 2), as reported in Table 10. Interestingly, $\mathcal{C_{\rm cut}}$ has a negligible effect on the bounds for dense graphs, where $\nu(G,\alpha)$ remains superior. Overall, this experiment suggests that the greedy edge clique cover provides a good trade-off between bound quality and computational time.
A notable exception is keller4, where $\mu(G,\mathcal{C}_{\rm cut})$ attains a tight upper bound of $12.64$ on $\alpha(G)=11$, computed in a reasonable time. To the best of the authors’ knowledge, this represents the strongest bound reported in the literature; see, e.g., Locatelli [41], Giandomenico et al. [30] and references therein.
7 Conclusions
We have presented a way to improve the classical Lovász $\theta$ bound for the stability number of a graph by applying the Lovász-Schrijver $N_{+}(\cdot)$ lift-and-project operator to tailored LP relaxations. Unlike previous approaches, some of the latter may not be constructed in polynomial time. In the case of clique inequalities, we have to handle exponentially many inequalities (with an associated NP-hard separation problem), while building strong nodal inequalities requires solving the SSP on specific subgraphs. We show that these expedients, if properly handled, are indeed helpful in letting the resulting SDP relaxations improve the $\theta$ bound without excessive extra effort. This study reveals that the Lovász and Schrijver $N_{+}(\cdot)$ operator can be of practical interest besides being a powerful theoretical tools. Therefore, its application to other combinatorial optimization problems that admit strong SDP relaxations, such as the graph coloring problem, is a natural research line. From a theoretical standpoint, a few questions remain open. First, from Theorem 3 we have ${N_{+}(\operatorname{NOD}(G,\Gamma))}\subseteq\operatorname{TH}^{+}(G)$, whereas the numerical experiments suggest ${N_{+}(\operatorname{NOD}(G,\Gamma))}=\operatorname{TH}^{+}(G)$. However, it still remains an open question. Secondly, the experiments showed a strong predominance of violated constraints lifted by multiplying by $x_{i}\geq 0$ on those multiplied by $1-x_{i}\geq 0$. This deserves a further theoretical investigation. Finally, an interesting research direction consists of finding classes of graphs where ${N_{+}(\operatorname{NOD}(G,\theta))}=\operatorname{STAB}(G)$, as for them the SSP would be polynomially solvable.
Acknowledgements
Conflict of interest
The authors declare that they have no conflict of interest.
Graph $\theta^{+}(G)$ $\mu(G,\mathcal{C_{\rm cut}})$ $\mu(G,\mathcal{C})$ $\nu(G,\alpha)$
Name UB % gap UB % gap UB % gap UB % gap
brock200_1 27.20 29.508 26.71 27.183 27.12 29.119 27.20 29.508
brock200_2 14.13 17.758 13.98 16.510 14.12 17.674 14.02 16.795
brock200_3 18.67 24.479 18.30 22.013 18.65 24.356 18.67 24.467
brock200_4 21.12 24.242 20.78 22.246 21.10 24.101 21.12 24.242
C125-9 37.55 10.431 35.87 5.487 36.23 6.566 37.55 10.430
DSJC125.1 38.04 11.896 36.42 7.125 36.79 8.207 38.04 11.881
DSJC125.5 11.40 14.021 11.04 10.359 11.36 13.598 11.35 13.531
keller4 13.47 22.417 12.64 14.873 13.39 21.709 13.45 22.236
p_hat300-1 10.02 25.253 9.81 22.622 10.02 25.194 8.58 7.288
p_hat300-2 26.71 6.855 26.00 4.015 26.49 5.950 26.69 6.768
p_hat300-3 40.70 13.056 39.74 10.384 40.44 12.321 40.70 13.053
p_hat500-1 13.01 44.533 12.95 43.847 13.01 44.526 10.68 18.721
p_hat500-2 38.56 48.306 37.94 45.934 38.38 47.597 38.54 48.225
p_hat700-1 15.05 36.774 15.02 36.589 15.05 36.773 11.30 2.709
p_hat700-2 48.44 10.091 47.89 8.831 48.24 9.647 48.41 10.023
sanr200_0.7 23.63 31.296 23.22 28.976 23.59 31.048 23.63 31.296
sanr200_0.9 48.90 16.439 47.49 13.061 48.03 14.356 48.90 16.440
sanr400_0.5 20.18 55.217 20.17 55.142 20.18 55.216 19.10 46.951
Table 9: Semidefinite bounds with improved edge clique covers.
Graph $\mu(G,\mathcal{C_{\rm cut}})$ $\mu(G,\mathcal{C})$ $\nu(G,\alpha)$
Name # iter Time # iter Time # iter Time
brock200_1 2.0 10.60 1.0 6.31 0.0 3.34
brock200_2 2.0 7.04 1.0 3.87 2.0 13.88
brock200_3 2.0 7.03 1.0 4.18 1.0 4.26
brock200_4 2.0 7.29 1.0 4.56 1.0 4.54
C125-9 4.0 32.91 4.0 33.90 1.0 9.95
DSJC125.1 4.0 32.81 4.0 33.64 1.0 9.22
DSJC125.5 2.0 5.53 1.0 2.97 1.0 3.64
keller4 4.0 29.15 1.0 11.57 1.0 15.01
p_hat300-1 2.0 36.83 1.0 21.78 3.0 262.60
p_hat300-2 3.0 311.08 2.0 168.35 1.0 232.25
p_hat300-3 4.0 75.37 2.0 43.34 1.0 36.93
p_hat500-1 1.0 37.54 1.0 40.53 5.0 854.09
p_hat500-2 3.0 841.79 2.0 588.09 1.0 927.59
p_hat700-1 1.0 76.46 1.0 78.54 7.0 6307.61
p_hat700-2 3.0 1405.20 2.0 1332.51 1.0 1208.12
sanr200_0.7 2.0 9.51 1.0 5.50 0.0 3.19
sanr200_0.9 3.0 43.39 3.0 43.01 1.0 20.16
sanr400_0.5 1.0 8.78 1.0 8.70 3.0 76.85
Table 10: CPU time and cutting-plane iterations with improved edge clique covers.